In the operating room, anesthesiologists use mass spectrometers tomonitor the respiratory gases of patients undergoing surgery. Onegas that is often monitored is the anesthetic isoflurane (molecularmass =
3.06 10-25 kg). In a spectrometer, a singly ionizedmolecule of isoflurane (charge = +e) moves at a speed of4.40 103 m/s on a circular path that has aradius of 0.105 m. What is the magnitudeof the magnetic field that the spectrometer uses?
1 T
Answer #1
The magnetic force is given by
           F= mv2/r
And F = Bqv
        Bqv =mv2/r
    The magnetic field that the spectrometeris
             B = mv/(qr)
   Here m=3.06 10-25 kg
           q = 1.6*10-19C
            v = 4.40 103m/s
            r = 0.105 m
Substitute the values.